# Elucidating the Design Space of Arbitrary-Noise-Based Diffusion Models

### v1

@@ 4.3 Experimental Results
The qualitative and quantitative results show that our EDA is more efficient and accurate in restoration. With only 5 sampling steps, EDA outperforms the 100-step Refusion and most task-specific methods, achieving SOTA performance in BFC and SR.
@@
Bias Field Correction in MRI. Quantitative (Tab. 1) and qualitative results (Fig. 3) demonstrate that our EDA effectively removes bias fields. The correlation coefficient (COCO) and coefficient of variation (CV) for both gray matter (GM) and white matter (WM) are additionally reported. EDA significantly enhances correction accuracy, as evidenced by the highest PSNR, SSIM, and COCO. Additionally, the lowest CV for WM and GM reflects that EDA has the most homogeneous intensity within the tissue. Notably, the restoration speed of EDA is $0.182$ sec/slice, significantly faster than the $9.665$ sec/slice of Refusion. Fig. 4 shows that EDA requires fewer sampling steps while yielding more accurate results. Fig. 3 shows the superior fidelity of our corrected images and the high accuracy of predicted bias fields (second row in Fig. 3).
@@
Metal Artifact Reduction in CT. Quantitative (Tab. 2) and qualitative results (Fig. 3) demonstrate that our EDA effectively reduces metal artifacts, exhibiting competitive performance and accurately restoring tissue structure. Notably, all MAR-specific methods use dual-domain (sinusoidal and image domain) information for restoration, whereas our EDA and Refusion rely only on image-domain information. Performance is thoroughly evaluated across A comprehensive evaluation was performed across five test data groups stratified by metal size. Using only the image domain, our EDA outperforms several dual-domain methods, showing that our EDA has strong image restoration performance, but a small performance gap remains between EDA and SOTA dual-domain methods. Additionally, the overall results of the EDA with 5-step sampling are better than the 100-step Refusion in Tab. 2, and the sampling pace and performance of EDA notably outperforms Refusion as shown in Fig. 4.
@@
Shadow Removal in Natural Images. Quantitative (Tab. 3) and qualitative (Fig. 3) results demonstrate that our EDA removes shadows most accurately and provides the most realistic restoration of non-shadow regions. The root mean square error (RMSE) in the LAB color space is reported additionally. Our method achieves the best performance of the overall image, as measured by PSNR, SSIM, and RMSE, which is also visually confirmed in Fig. 3. The EDA restoration model is based on ShadowFormer, and EDA achieves improved performance, highlighting its effectiveness. Our EDA significantly enhances both the sampling efficiency and the de-shading effect compared to Refusion, as shown in Fig. 4.
@@ 5 Conclusion
This paper introduces EDA, a unified design space for diffusion models with arbitrary noise patterns for universal image transformation. EDA generalizes vanilla design space to multivariate Gaussian distributions and rigorously demonstrates its ability to diffuse arbitrary noise patterns. Furthermore, our study proves that extending the diffused noise from simple Gaussian patterns, as used in EDM, to more complex arbitrary patterns introduces no additional computational overhead. EDA has been validated on MRI bias field correction, CT metal artifact reduction, and natural image shadow removal, achieving superior accuracy and efficiency in fewer than five steps as shown in Fig. 4.
@@ Appendix A Derivation of $f(\cdot)$, $g(\cdot)$ and $\phi(\cdot)$ in SDE
$P\left(x_{t}\mid x_{0}\right)=\mathcal{N}\left(x_{t};\,s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\,\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\right),$ (21)
@@ 1. Derivation of the Mean Component
$\frac{\mathrm{d}\mu(t)}{\mathrm{d}t}=f(t)\mu(t)+\phi_{x_{0}}(t),\quad\mu(0)=x_{0}$ (28)
@@
$\mu(t)=\exp\left(\int_{0}^{t}f(\tau)\mathrm{d}\tau\right)x_{0}+\exp\left(\int_{0}^{t}f(\tau)\mathrm{d}\tau\right)\int_{0}^{t}\phi(\tau)\exp\left(-\int_{0}^{\tau}f(s)\mathrm{d}s\right)\mathrm{d}\tau$
@@ 2. Derivation of the Variance Component
$\Sigma(t)=\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}},\quad\text{where }\Sigma_{x_{0}}=\sum_{m=1}^{M}h_{m,\bm{x}_{0}}h_{m,\bm{x}_{0}}^{\top}$ (31)
@@
$\frac{\mathrm{d}\Sigma_{S}(t)}{\mathrm{d}t}=2f(t)\Sigma_{S}(t)+g^{2}(t)\sum_{m=1}^{M}h_{m,\bm{x}_{0}}h_{m,\bm{x}_{0}}^{\top},\quad\Sigma(0)=0$ (32)
@@ Appendix B Derivation of Probability Flow ODE for EDA
[…]
@@ Appendix C Derivation of Deterministic Sampling Formula from PFODE
$f(t)=\frac{s^{\prime}(t)}{s(t)},\quad\phi_{x_{0}}(t)=\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\quad g^{2}(t)=\frac{2s^{2}(t)\sigma(t)\sigma^{\prime}(t)}{(\eta+1)^{2}}$ (50)
@@
$\frac{\mathrm{d}x}{\mathrm{d}t}=\frac{s^{\prime}(t)}{s(t)}x+\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-\frac{s^{2}(t)\sigma(t)\sigma^{\prime}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\nabla_{x_{t}}\log p(x|x_{0})$ (51)
@@
$=\int_{\mathbb{R}^{d}}\mathcal{N}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\right)\|D_{x_{0}}(x;\sigma)-x_{0}\|^{2}\mathrm{d}x$ (54)
$\triangleq\int_{\mathbb{R}^{d}}\mathcal{L}_{x_{0}}(D;\sigma(t),x)\mathrm{d}x$ (55)
@@
00 0 $=\nabla_{D_{x_{0}}(x,\sigma)}\mathcal{L}_{x_{0}}(D;\sigma(t),x)$ (56)
00 0 $=\nabla_{D_{x_{0}}(x,\sigma)}\left[\mathcal{N}\left(s(t)x_{0}+\frac{s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}\eta h_{m,\bm{x}_{0}},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\right)\|D_{x_{0}}(x;\sigma)-x_{0}\|^{2}\right]$ (57)
00 0 $=\nabla_{D_{x_{0}}(x,\sigma)}\|D_{x_{0}}(x;\sigma)-x_{0}\|^{2}$ (58)
@@
For the conditional distribution:
$p(x_{t}|x_{0})=\mathcal{N}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\right)$ (60)
Then derive the score function:
$\nabla_{x_{t}}\log p(x_{t}|x_{0})$ $=\dfrac{\nabla_{x_{t}}p(x_{t}|x_{0})}{p(x_{t}|x_{0})}$ (61)
$=\dfrac{\nabla_{x_{t}}\mathcal{N}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\right)}{\mathcal{N}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\right)}$ (62)
where the numerator is:
$\nabla_{x_{t}}\mathcal{N}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\right)$ (64)
$=\mathcal{N}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\right)$
$\quad\nabla_{x_{t}}\left[-\frac{(\eta+1)^{2}}{2s^{2}(t)\sigma^{2}(t)}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-x_{t}\right)^{\top}\Sigma_{x_{0}}^{-1}\right.$
$\qquad\left.\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-x_{t}\right)\right]$ (65)
$=\mathcal{N}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\right)$
$\quad\left[\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\Sigma_{x_{0}}^{-1}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-x_{t}\right)\right]$ (66)
@@
$\nabla_{x_{t}}\log p(x_{t}|x_{0})$ $=\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\Sigma_{x_{0}}^{-1}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-x_{t}\right)$ (67)
@@
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=\frac{s^{\prime}(t)}{s(t)}x+\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-\frac{s^{2}(t)\sigma(t)\sigma^{\prime}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\left(\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\Sigma_{x_{0}}^{-1}\right.$
@@
$=\frac{s^{\prime}(t)}{s(t)}x+\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-\frac{\sigma^{\prime}(t)}{\sigma(t)}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-x\right)$ (70)
@@
$\mathcal{L}(D;\sigma(t))$ $=\mathbb{E}_{y\sim P_{\text{data}}}\mathbb{E}_{x\sim\mathcal{N}(\mu(y,t),\bar{\Sigma}(t))}\|D(x;\sigma)-D_{y}(x;\sigma)\|^{2}$ (74)

### v2

@@ 4.3 Experimental Results
The qualitative and quantitative results show that our EDA is more efficient and accurate in restoration. With only 5 sampling steps, EDA outperforms the 100-step Refusion and most task-specific methods, achieving SOTA performance in BFC and SR. The MeanFlow training objective fundamentally struggles to align and minimize the complex Task-Specific Error Distribution characteristic of inverse problems. This results in a significant inherent bias in the MeanFlow prediction, leading to sub-optimal performance compared to SOTA restoration methods in the PSNR/SSIM-based precise recovery tasks. Consequently, its scores are excluded from the qualitative analysis tables.
@@
Bias Field Correction in MRI. Quantitative (Tab. 1) and qualitative results (Fig. 3) demonstrate that our EDA effectively removes bias fields. The correlation coefficient (COCO) and coefficient of variation (CV) for both gray matter (GM) and white matter (WM) are additionally reported. EDA significantly enhances correction accuracy, as evidenced by the highest PSNR, SSIM, and COCO. Additionally, the lowest CV for WM and GM reflects that EDA has the most homogeneous intensity within the tissue, ensuring the homogeneity required for precise subsequent tissue segmentation. MeanFlow exhibits the highest CV for GM ($15.49$) among all methods, indicating that its ODE training process struggles to achieve the necessary tissue intensity uniformity. Notably, the restoration speed of EDA is $0.182$ sec/slice, significantly faster than the $9.665$ sec/slice of Refusion, representing a substantial $\sim 53$-fold acceleration, enabling EDA viable for high-throughput clinical applications. Fig. 4 shows that EDA requires fewer sampling steps while yielding more accurate results. Fig. 3 shows the superior fidelity of our corrected images and the high accuracy of predicted bias fields (second row in Fig. 3).
@@
Metal Artifact Reduction in CT. Quantitative (Tab. 2) and qualitative results (Fig. 3) demonstrate that our EDA effectively reduces metal artifacts, exhibiting competitive performance and accurately restoring tissue structure. Notably, all MAR-specific methods use dual-domain (sinusoidal and image domain) information for restoration, whereas our EDA and Refusion rely only on image-domain information. A comprehensive evaluation was performed across five test data groups stratified by metal size. Using only the image domain, our EDA outperforms several dual-domain methods, showing that our EDA has strong image restoration performance, but a small performance gap remains between EDA and SOTA dual-domain methods. MeanFlow exhibits low average performance because the task involves severe data loss and multiple plausible restorations, but ODE diffusion produces an averaged solution, resulting in blurring in artifact regions. Additionally, the overall results of the EDA with 5-step sampling are better than the 100-step Refusion in Tab. 2, and the sampling pace and performance of EDA notably outperforms Refusion as shown in Fig. 4.
@@
Shadow Removal in Natural Images. Quantitative (Tab. 3) and qualitative (Fig. 3) results demonstrate that our EDA removes shadows most accurately and provides the most realistic restoration of non-shadow regions. The root mean square error (RMSE) in the LAB color space is reported additionally. Our method achieves the best performance of the overall image, as measured by PSNR, SSIM, and RMSE, which is also visually confirmed in Fig. 3. The EDA restoration model is based on ShadowFormer, and EDA achieves improved performance, highlighting its effectiveness. Crucially, EDA maintains the highest fidelity in the Non-Shadow (NS) region (PSNR $34.31$ dB, RMSE $3.77$), demonstrating that the generalized diffusion framework preserves original image quality better than all competitors, indicating EDA’s ability to precisely delineate the shadow boundary and avoid introducing artifacts into the background. Our SDE diffusion framework significantly outperforms ODE MeanFlow on this task. Shadow removal is ill-posed problem that requires the model to generate plausible, high-frequency textures and colors that were occluded by the shadow [25]. MeanFlow’s ODE training follows an averaged trajectory, which leads to a high ALL RMSE of $9.77$. In contrast, the SDE-based stochasticity integrated by EDA explores the manifold of plausible restorations and corrects errors, which is critical for finding a high-fidelity solution. Our EDA significantly enhances both the sampling efficiency and the de-shading effect compared to Refusion, as shown in Fig. 4.
@@ 5 Conclusion
This paper introduces EDA, a unified design space for diffusion models with arbitrary noise patterns for universal image restoration. EDA generalizes vanilla design space to multivariate Gaussian distributions and rigorously demonstrates its ability to diffuse arbitrary noise patterns. Furthermore, our study proves that extending the diffused noise from simple Gaussian patterns, as used in EDM, to more complex arbitrary patterns introduces no additional computational overhead. EDA has been validated on MRI bias field correction, CT metal artifact reduction, and natural image shadow removal, achieving superior accuracy and efficiency in fewer than $5$ steps as shown in Fig. 4.
Limitations: Compared to deterministic diffusion methods, such as Flow Matching (ODE diffusion) and Cold Diffusion (operator-based), EDA employs an SDE-driven diffusion process where inherent stochasticity enhances robustness and performance. While Prop. 1 presents three configurations for all potential scenarios, EDM provides a smooth transition that balances stochasticity and flexibility. Case 1 maximizes randomness by theoretically spanning the entire noise domain via weighted basis functions. In contrast, the more widely applicable Cases 2–3 reduce randomness through only stochastic scalar weight adjustments, resembling deterministic methods. Consequently, EDA’s design entails an inherent trade-off between diffusion stochasticity and application scope.
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant No. 82527807; the Shenzhen Medical Research Fund under Grant No. C2501016; the Science and Technology Innovation Committee of Shenzhen Municipality under Grant No. JCYJ20250604145426036; the Key Research & Development Program of Heilongjiang Province under Grant Nos. 2023ZX01A08;
Supplementary Material
@@ Appendix A Derivation of $f(\cdot)$, $g(\cdot)$ and $\phi(\cdot)$ in SDE
$P(x_{t}\mid x_{0})$ $=\mathcal{N}\Bigl(x_{t};\,s(t)x_{0}+\frac{\eta\,s(t)\,\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},$ (22)
$\hskip 86.40024pt\frac{s^{2}(t)\,\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\Bigr).$
@@ 1. Derivation of the Mean Component
$\frac{\mathrm{d}\mu(t)}{\mathrm{d}t}=f(t)\mu(t)+\phi_{x_{0}}(t),\qquad\mu(0)=x_{0}$ (29)
@@
$\mu(t)=\exp\!\Bigl(\!\int_{0}^{t}f(\tau)\,\mathrm{d}\tau\!\Bigr)x_{0}$
$+\exp\!\Bigl(\!\int_{0}^{t}f(\tau)\,\mathrm{d}\tau\!\Bigr)\int_{0}^{t}\phi(\tau)\exp\!\Bigl(\!-\!\int_{0}^{\tau}f(s)\,\mathrm{d}s\!\Bigr)\,\mathrm{d}\tau.$
@@ 2. Derivation of the Variance Component
$\Sigma(t)=\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}},\qquad$ (32)
where $\Sigma_{x_{0}}=\sum_{m=1}^{M}h_{m,\bm{x}_{0}}h_{m,\bm{x}_{0}}^{\top}$.
@@
$\small\frac{\mathrm{d}\Sigma_{S}(t)}{\mathrm{d}t}=2f(t)\Sigma_{S}(t)+g^{2}(t)\sum_{m=1}^{M}h_{m,\bm{x}_{0}}h_{m,\bm{x}_{0}}^{\top},\quad\Sigma(0)=0$ (33)
@@ Appendix B Derivation of Probability Flow ODE for EDA
where $\nabla\cdot$ denotes divergence and $\nabla\nabla:$ denotes the double divergence operator.
@@ Appendix C Derivation of Deterministic Sampling Formula from PFODE
$f(t)$ $=\frac{s^{\prime}(t)}{s(t)},\quad\phi_{x_{0}}(t)=\frac{\eta\,s(t)\,\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},$
$g^{2}(t)$ $=\frac{2\,s^{2}(t)\,\sigma(t)\,\sigma^{\prime}(t)}{(\eta+1)^{2}}.$ (51)
@@
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=\frac{s^{\prime}(t)}{s(t)}\,x+\frac{\eta\,s(t)\,\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}$ (52)
$\quad-\frac{s^{2}(t)\,\sigma(t)\,\sigma^{\prime}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\,\nabla_{x_{t}}\!\log p(x\mid x_{0}).$
@@
$\mathcal{L}_{x_{0}}(D;\sigma(t))=\mathbb{E}_{x\sim\mathcal{N}(\mu(x_{0},t),\bar{\Sigma}(t))}\|D(x;\sigma)-x_{0}\|^{2},$ (54)
$={\small\int_{\mathbb{R}^{d}}\mathcal{N}\!\Bigl(s(t)x_{0}+\tfrac{\eta\,s(t)\,\sigma(t)}{\eta+1}\!\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\;\tfrac{s^{2}(t)\,\sigma^{2}(t)}{(\eta+1)^{2}}\,\Sigma_{x_{0}}\Bigr)\,}$
$\qquad{}\times\|D_{x_{0}}(x;\sigma)-x_{0}\|^{2}\,\mathrm{d}x,$ (55)
$\triangleq\int_{\mathbb{R}^{d}}\mathcal{L}_{x_{0}}(D;\sigma(t),x)\,\mathrm{d}x.$ (56)
@@
$0$ $=\nabla_{D_{x_{0}}(x,\sigma)}\mathcal{L}_{x_{0}}(D;\sigma(t),x),$ (57)
$0$ $=\small\begin{aligned} \nabla_{D_{x_{0}}(x,\sigma)}\!\Bigl[\mathcal{N}\!\Bigl(s(t)x_{0}+\tfrac{s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}\eta h_{m,\bm{x}_{0}},\;\\ \tfrac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\Bigr)\hskip 18.49988pt{}\times\|D_{x_{0}}(x;\sigma)-x_{0}\|^{2}\Bigr]\end{aligned}$ (58)
$0$ $=\nabla_{D_{x_{0}}(x,\sigma)}\|D_{x_{0}}(x;\sigma)-x_{0}\|^{2}.$ (59)
@@
For the conditional distribution, we define the mean $\bm{\mu}_{c}$ and covariance $\bm{\Sigma}_{c}$ as:
$\bm{\mu}_{c}=s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\qquad\bm{\Sigma}_{c}=\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}.$ (61)
The distribution is then expressed compactly as $p(x_{t}|x_{0})=\mathcal{N}(x_{t};\bm{\mu}_{c},\bm{\Sigma}_{c})$. The score function is derived as:
$\nabla_{x_{t}}\log p(x_{t}|x_{0})$ $=\frac{\nabla_{x_{t}}\mathcal{N}(x_{t};\bm{\mu}_{c},\bm{\Sigma}_{c})}{\mathcal{N}(x_{t};\bm{\mu}_{c},\bm{\Sigma}_{c})}$ (62)
$=\frac{\mathcal{N}(x_{t};\bm{\mu}_{c},\bm{\Sigma}_{c})\left[\bm{\Sigma}_{c}^{-1}(\bm{\mu}_{c}-x_{t})\right]}{\mathcal{N}(x_{t};\bm{\mu}_{c},\bm{\Sigma}_{c})}$ (63)
$=\bm{\Sigma}_{c}^{-1}(\bm{\mu}_{c}-x_{t}).$ (64)
Substituting the definitions of $\bm{\mu}_{c}$ and $\bm{\Sigma}_{c}$ back yields the final analytical form.
@@
$\nabla_{x_{t}}\log p(x_{t}|x_{0})$ $=\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\Sigma_{x_{0}}^{-1}\Bigl(s(t)x_{0}$
$+\frac{\eta\,s(t)\,\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-x_{t}\Bigr).$ (65)
@@
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=\frac{s^{\prime}(t)}{s(t)}x+\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}$
$-\frac{s^{2}(t)\sigma(t)\sigma^{\prime}(t)}{(\eta+1)^{2}}\Sigma_{x_{0}}\left(\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\Sigma_{x_{0}}^{-1}\right$.
@@
$=\frac{s^{\prime}(t)}{s(t)}x+\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}$
$-\frac{\sigma^{\prime}(t)}{\sigma(t)}\left(s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-x\right)$ (67)
@@
$\mathcal{L}(D;\sigma(t))$
$=\mathbb{E}_{y\sim P_{\text{data}}}\mathbb{E}_{x\sim\mathcal{N}(\mu(y,t),\bar{\Sigma}(t))}\|D(x;\sigma)-D_{y}(x;\sigma)\|^{2}$ (71)